# MUSE-FM: Multi-task Environment-aware Foundation Model for Wireless Communications

### v1

@@ Abstract
Recent advancements in foundation models (FMs) have attracted increasing attention in the wireless communication domain. Leveraging the powerful multi-task learning capability, FMs hold the promise of unifying multiple tasks of wireless communication with a single framework. Nevertheless, existing wireless FMs face limitations in the uniformity to address multiple tasks with diverse inputs/outputs across different communication scenarios. In this paper, we propose a MUlti-taSk Environment-aware FM (MUSE-FM) with a unified architecture to handle multiple tasks in wireless communications, while effectively incorporating scenario information. Specifically, to achieve task uniformity, we propose a unified prompt-guided data encoder-decoder pair to handle data with heterogeneous formats and distributions across different tasks. Besides, we integrate the environmental context as a multi-modal input, which serves as prior knowledge of environment and channel distributions and facilitates cross-scenario feature extraction. Simulation results illustrate that the proposed MUSE-FM outperforms existing methods for various tasks, and its prompt-guided encoder-decoder pair improves the scalability for new task configurations. Moreover, the incorporation of environment information improves the ability to adapt to different scenarios.
Index Terms:
foundation models (FMs), multi-task learning, scene graph, environment awareness
@@ I Introduction
The integration of artificial intelligence (AI) with wireless communication has become increasingly vital for the design and optimization of sixth-generation (6G) wireless communication networks [1, 2]. Specifically, the integration of AI demonstrates performance improvement or complexity reduction for various communication tasks, including channel estimation and prediction, channel state information (CSI) feedback, channel decoding, and resource allocation under dynamic conditions [3]. The recent paradigm shift towards Foundation Models (FMs), characterized by a large amount of parameters and pretraining on vast datasets, has dramatically reshaped the AI landscape [4]. These models exhibit remarkable reasoning capabilities and generalizability across multiple domains and multiple tasks. Given their profound impact across diverse domains, exploring the applicability and potential of FMs in 6G wireless communications emerges as a crucial research imperative to unlock next-generation wireless intelligence.
@@ I-A Prior Works
There is an emerging trend of leveraging FMs to support massive device density, high-data-rate communications, and ubiquitous intelligence in 6G networks [5, 6, 7]. To be specific, existing works mainly have attempted to harness three key advantages of FMs, namely powerful feature extraction capability, few-shot learning capability, and generalization capability, to address challenges in wireless communication systems. Firstly, owing to the substantial parameters, FMs acquire superior feature extraction capabilities compared to traditional DL architectures. This inherent advantage has attracted a growing research focus on leveraging these models to address complex problems and scenarios for various tasks in wireless communications, including channel prediction and scatter generation, etc. Specifically, Liu et al. [8] first introduce FM to physical layer communications, and propose LLM4CP to deal with the challenge of channel prediction in high-dynamic scenarios. LLM4CP unleashes the power of FMs to achieve accurate channel prediction. It is proposed by Li et al. in [9] that an FM-enabled decomposition-based multi-objective evolutionary algorithm can jointly optimize the communication and sensing objectives in integrated sensing and communication (ISAC) systems. Han et al. [10] focus on the scatterer generation task from LiDAR point clouds in multiple scenarios with multiple frequency bands, which demonstrates the advantages of FM in handling complex tasks and adapting to dynamic environments. Besides, FMs are also employed in fluid antenna systems [11] to tackle the challenges associated with user mobility.
@@
Secondly, pretrained on extensive datasets, FMs possess inherent few-shot/zero-shot learning capabilities, compared to conventional deep learning (DL)-based models. The artificial general intelligence acquired by the FMs through pretraining enables direct task inference without the need to design and train dedicated networks from scratch. For example, in [12], the authors adopt a pretrained FM to perform power allocation using a few-shot learning approach. In this method, the channel gains and the corresponding transmit power strategies are contacted as prompts and fed to the FM. Then, the FM can provide the power allocation strategy without any fine-tuning. Thirdly, FMs exhibit better generalization capability to different wireless environments, compared to traditional DL models. For instance, leveraging this generalization capability, Guo et al. [13] propose an FM-based CSI feedback method that is adaptable to various scenarios. The authors incorporate the channel distribution in different environments as input and enable the model to handle multiple scenarios simultaneously.
@@
It can be observed that FMs successfully improve the performance and robustness in various wireless communication tasks [8, 9, 10, 11, 12, 13]. Nevertheless, existing works typically design a dedicated FM for each individual task. Consequently, developing separate FMs for all the aforementioned tasks would incur substantial storage and deployment overhead. Thus, recently, a few initial attempts have been made to propose a multi-task FM to unify multiple tasks with a single model. In [14], the authors propose an FM to extract features of wireless channels, which are used as common inputs of downstream models in downstream tasks.  [15, 16] also focus on channel-associated tasks such as channel estimation, beam selection and path loss estimation. They employ an FM to jointly learn a unified representation for multiple tasks, with task-specific adapters or fine-tuning for different downstream tasks. The authors in [17] attempt to integrate tasks with different data formats and distributions to fully demonstrate the multitasking capability of FMs. For this purpose, the authors propose a multi-task FM to unify three tasks, namely channel prediction, signal detection and multi-user precoding, with the aid of task-specific adapters and elaborately designed prompts.
@@ I-B Motivations
Limitation in task uniformity: Existing works suffer from the narrow task scope, as they predominantly focus on channel-related tasks. These works utilize pilot signals or CSI as input, employing FMs to extract features for multiple downstream tasks. However, they fail to unify other critical components across the entire physical layer functionalities. Besides, existing works employ task-specific data encoders and decoders to handle diverse input/output formats and feature spaces across different tasks and parameters. However, given the diverse tasks and parameter variations, maintaining separate data encoder-decoder pairs for each task and parameter still incurs high computational and deployment overhead. It also degrades the scalability of the model, contradicting the objectives of unified multi-task FM paradigms.
Limitation in scenario uniformity: Scenario uniformity denotes the environmental adaptability of the model to generalize across multiple scenarios, which is critical for real-world deployment. However, existing multi-task FMs are typically optimized for a specific scene. When encountered a new scenario with a different spatial structure, the shift in distribution of channel characteristics leads to severe performance distortion. In this case, data samples in the new environment are required to be collected and online fine-tuning is then performed, which imposes substantial computational overheads especially for FMs with large-scale parameters. Therefore, existing works significantly compromise the environmental adaptability of FMs, and increase deployment costs.
@@ I-C Our Contributions
  • In addition, to improve the environmental uniformity of the proposed method, we integrate the environmental context as a multi-modal input. Specifically, prior knowledge of environment and channel distributions is obtained from scene graphs and facilitates cross-scenario feature extraction. Thus, we establish a multi-task environment-aware FM that dynamically adapts to varying propagation environments.
@@ II-A The Structure of MIMO-OFDM Transceivers
The typical structure of a MIMO-OFDM transceiver, deployed at the BS, is depicted in Fig. 1. The MIMO-OFDM transmitter begins with channel encoding (e.g., LDPC or polar codes), which adds redundancy to the input bitstream $\mathbf{b}$ for robustness against channel errors. The encoded bits $\mathbf{s}$ are then mapped to complex symbols by modulation (e.g., QAM or PSK). For multi-user systems, multi-user precoding is applied to minimize inter-user interference and simultaneously transmit $\mathbf{x}_{k},k=1,2,\cdots,K$ for $K$ users, based on the multi-user CSI. The modulated symbols are subsequently allocated to different OFDM subcarriers, added by a cyclic prefix (CP) to mitigate inter-symbol interference (ISI) caused by multi-path fading. The data transmitted to the $k$-th user of the $m$-th subcarrier is referred to as ${x}_{m,k},k=1,2,\cdots,K,m=1,2,\cdots,M$.
Due to the sparse scattering environment of the mmWave frequency [18], we adopt the Saleh-Valenzuela channel model [19] for the multi-path channel $\mathbf{h}_{m,k}\in\mathbb{C}^{N_{t}\times 1}$ between the BS and the user $k$ at the $m$-th subcarrier as
$\mathbf{h}_{m,k}=\sum_{l=1}^{L_{k}}\beta_{m,k,l}e^{-j2\pi f_{m}\tau_{k,l}}\mathbf{a}_{m}(\theta_{k,l},\phi_{k,l}),$ (1)
@@
$\mathbf{a}_{m}(\theta,\phi)$ $=\frac{1}{\sqrt{N_{t}}}[1,\cdots,e^{jk_{m}d(n_{v}\sin\phi\sin\theta+n_{h}\cos\theta)},\cdots,$
$e^{jk_{m}d((N_{v}-1)\sin\phi\sin\theta+(N_{h}-1)\cos\theta)}]^{T},$ (2)
@@
Building upon the aforementioned description of the MIMO-OFDM transceiver, we select critical yet challenging tasks including channel estimation, MIMO detection, channel decoding, and multi-user precoding, while ingoring the relatively trivial components (e.g., channel encoding, which requires only a matrix multiplication operation). The task selection aims to establish a unified FM capable of addressing the most demanding aspects of a transceiver. Furthermore, we incorporate user localization as an essential task, given its fundamental role in 6G ISAC systems, a pivotal direction for next-generation wireless networks. In the next subsection, we will formulate the selected tasks individually.
@@ III-A Overall Framework
The framework of the proposed model is presented in Fig. 2, which includes the scene encoder, prompted-guided unified encoder-decoder pair, and FM backbone. To address scenario limitation, leveraging the strong correlation between physical environments and channel characteristics [21], we attempt to incorporate environmental information to allow cross-scenario generalization in multi-task FM. Specifically, the spatial structure of the physical environment, including the area building distribution as well as the BS position, determines the propagation path information of the wireless channel, and thus directly impacts channel characteristics. Therefore, involving the environmental knowledge as a multi-modal input provides a promising way to inform FMs of the channel distribution and improves the generalization ability in different scenarios. We implement it through a Vision Transformer-based encoder that extracts environmental and channel prior knowledge from scene graphs. The prior knowledge is subsequently fused with wireless data to enhance the model’s environmental adaptability. The specific design of the scene encoder is described in Section III-B.
@@
To handle data with heterogeneous formats and distributions across multiple tasks, we propose a prompt-guided unified data encoder whose parameters are dynamically generated by task-specific instructions through hypernetworks. In particular, the task instruction is fed to a hypernetwork to dynamically generate parameters, which is utilized in the data encoder. This architecture allows the instruction to “guide” the encoder to adaptively extract task-specific features according to each task’s objectives and data characteristics. To achieve this, two key components are included: a parameter generator, i.e., which is used to generate task-specific parameters from task instruction via a text encoder; a prompted-guided unified encoder-decoder pair, to utilize the generated parameters to project the data of multiple tasks into the shared FM space (or generate the final outputs in multiple tasks). The two components are elaborated respectively in Section III-C and III-D.
@@
The token embeddings from scene graphs and wireless data (e.g., CSI and received symbols) are concatenated to form the input embeddings for the FM backbone. The FM backbone consists of position embedding and stacked classical transformer blocks, which is illustrated in Section III-E. Equipped with large-scale parameters and a layered structure, the FM backbone facilitates complex feature assimilation, thereby enhancing its expressive capacity for multiple tasks. Finally, similar to the data encoder, equipped with task-aware parameters, the prompt-guided unified decoder generates the output for different tasks.
@@ III-B Scene Encoder
Figure 3: Vision-transformer based scene encoder.
The structure of the scene encoder to effectively exploit environmental information is illustrated in Fig. 3. Direct utilization of raw 3D environmental data is challenging due to its high dimensionality. We therefore implement preprocessing, transforming environmental information into compact representations suitable for encoder input. Given that top-down views holistically encapsulate the spatial structure of the environment, that is, scene layouts and obstacle distributions, we extract it to construct the scene graph [21]. This scene graph is then discretized into a matrix form, where the value 1 represents the obstacles, and the value 0 represents free space.
@@ III-D1 Pre-processor and unified data encoder
In Fig. 4, we take channel estimation as an example to illustrate how the original data input is preprocessed and handled by the unified data encoder. The data embeddings $\mathrm{X}_{n}^{\rm emb}$ are concatenated with the scene embeddings $\mathrm{X}_{s}^{\rm emb}$. Besides, we use the standard approach of adding an extra learnable “classification token” (cls token) to the sequence, to form the final input of the FM backbone. We refer to the input as $\mathrm{X}_{\rm FM}^{\rm emb}$.

### v2

@@ Abstract
Recent advancements in foundation models (FMs) have attracted increasing attention in the wireless communication domain. Leveraging the powerful multi-task learning capability, FMs hold the promise of unifying multiple tasks of wireless communication with a single framework. Nevertheless, existing wireless FMs face limitations in the uniformity to address multiple tasks with diverse inputs/outputs across different communication scenarios. In this paper, we propose a MUlti-taSk Environment-aware FM (MUSE-FM) with a unified architecture to handle multiple tasks in wireless communications, while effectively incorporating scenario information. Specifically, to achieve task uniformity, we propose a unified prompt-guided data encoder-decoder pair to handle data with heterogeneous formats and distributions across different tasks. Besides, we integrate the environmental context as a multi-modal input, which serves as prior knowledge of environment and channel distributions and facilitates cross-scenario feature extraction. Simulation results illustrate that the proposed MUSE-FM outperforms existing methods for various tasks, and its prompt-guided encoder-decoder pair facilitates few-shot adaptation to new task configurations. Moreover, the incorporation of environment information improves the ability to adapt to different scenarios.
@@ I Introduction
The integration of artificial intelligence (AI) with wireless communication has become increasingly vital for the design and optimization of sixth-generation (6G) wireless communication networks [1, 2]. Specifically, the integration of AI demonstrates performance improvement or complexity reduction for various communication tasks, including channel estimation and prediction, channel state information (CSI) feedback, channel decoding, and resource allocation under dynamic conditions [3]. The recent paradigm shift towards Foundation Models (FMs), characterized by a large amount of parameters and pretraining on vast datasets, has dramatically reshaped the AI landscape [4]. These models exhibit remarkable reasoning capabilities and generalizability across multiple domains and tasks. Given their profound impact across diverse domains, exploring the applicability and potential of FMs in 6G wireless communications emerges as a crucial research imperative to unlock next-generation wireless intelligence.
@@ I-A Prior Works
There is an emerging trend of leveraging FMs to support massive device density, high-data-rate communications, and ubiquitous intelligence in 6G networks [5, 6, 7]. To be specific, existing works mainly have attempted to harness three key advantages of FMs, namely powerful feature extraction, few-shot learning, and generalization capability, to address challenges in wireless communication systems. Firstly, owing to the substantial parameters, FMs acquire superior feature extraction capabilities compared to traditional deep learning (DL) architectures. This inherent advantage has attracted a growing research focus on leveraging these models to address complex problems and scenarios for various tasks in wireless communications, such as channel prediction and scatter generation. Specifically, Liu et al. [8] introduce FM to physical layer communications, and propose LLM4CP to deal with the challenge of channel prediction in high-dynamic scenarios. LLM4CP unleashes the power of FMs to achieve accurate channel prediction. It is proposed in [9] that an FM-enabled decomposition-based multi-objective evolutionary algorithm can jointly optimize the communication and sensing objectives in integrated sensing and communication (ISAC) systems. [10] focuses on beam prediction task for near-field communications, which demonstrates the advantages of FM in handling complex temporal dynamics due to the joint dependence on user angle and distance for near field channels. Besides, FMs are also employed in fluid antenna systems [11] to tackle the challenges of user mobility.
@@
Secondly, pretrained on extensive datasets, FMs possess inherent few-shot/zero-shot learning capabilities, compared to conventional DL-based models. The artificial general intelligence acquired by the FMs through pretraining enables direct task inference without the need to design and train dedicated networks from scratch. For example, in [12], the authors adopt a pretrained FM to perform power allocation using a few-shot learning approach. In this method, the channel gains and the corresponding transmit power strategies are contacted as prompts and fed to the FM. Then, the FM can provide the power allocation strategy without any fine-tuning. Thirdly, FMs exhibit better generalization capability to different wireless environments, compared to traditional DL models. For instance, leveraging this generalization capability, Guo et al. [13] propose an FM-based CSI feedback method that is adaptable to various scenarios. The authors incorporate the channel distribution in different environments as input and enable the model to handle multiple scenarios simultaneously.
@@
It has been observed that FMs have successfully improved the performance and robustness in various wireless communication tasks [8, 9, 10, 11, 12, 13]. Nevertheless, existing works typically design a dedicated FM for each individual task. Consequently, developing separate FMs for all the aforementioned tasks would incur substantial storage and deployment overhead. Recently, a few initial attempts have been made to propose a multi-task FM to unify multiple tasks with a single model. In [14], the authors propose an FM to extract features of wireless channels, which are used as common inputs of downstream models in downstream tasks. There are also studies  [15, 16] focusing on channel-associated tasks such as channel estimation, beam selection and path loss estimation. They employ an FM to jointly learn a unified representation for multiple tasks, with task-specific adapters or fine-tuning for different downstream tasks. The authors in [17] attempt to integrate tasks with different data formats and distributions to fully demonstrate the multitasking capability of FMs. For this purpose, a multi-task FM is proposed to unify three tasks, namely channel prediction, signal detection and multi-user precoding, with the aid of task-specific adapters and elaborately designed prompts.
@@ I-B Motivations
Limitation in task uniformity: Existing works suffer from the narrow task scope, as they predominantly focus on channel-related tasks. These works utilize pilot signals or CSI as input, employing FMs to extract features for multiple downstream tasks. However, they fail to unify other critical components across the entire physical layer functionalities. Besides, existing works employ task-specific data encoders and decoders to handle diverse input/output formats and feature spaces across different tasks and parameters. However, given the diverse tasks and parameter variations, maintaining separate data encoder-decoder pairs for each task and parameter incurs high computational and deployment overhead. It also degrades the scalability of the model, contradicting the objectives of unified multi-task FM paradigms.
Limitation in scenario uniformity: Scenario uniformity refers to the environmental adaptability of the model to generalize across multiple scenarios, which is critical for real-world deployment. However, existing multi-task FMs are typically optimized for a specific scene. When encountered a new scenario with a different spatial structure, the shift in distribution of channel characteristics leads to severe performance distortion. In this case, data samples in the new environment are required to be collected and online fine-tuning is then performed, which imposes substantial computational overheads especially for FMs with large-scale parameters. Therefore, existing works significantly compromise the environmental adaptability of FMs, and increase deployment costs.
@@ I-C Our Contributions
  • In addition, to improve the environmental uniformity, we integrate the environmental context as a multi-modal input. Specifically, prior knowledge of environment and channel distributions is obtained from scene graphs and facilitates cross-scenario feature extraction. Thus, we establish a multi-task environment-aware FM that dynamically adapts to varying propagation environments.
@@ II-A The Structure of MIMO-OFDM Transceivers
The typical structure of a MIMO-OFDM transceiver, deployed at the BS, is depicted in Fig. 1. The MIMO-OFDM transmitter begins with channel encoding (e.g., LDPC or polar codes), which adds redundancy to the input bitstream $\mathbf{b}$ for robustness against channel errors. The encoded bits $\mathbf{s}$ are then mapped to complex symbols by modulation (e.g., QAM or PSK). For multi-user systems, multi-user precoding is applied to minimize inter-user interference and simultaneously transmit $\mathbf{x}_{k},k=1,2,\cdots,K$ for $K$ users, based on the multi-user CSI. The modulated symbols are subsequently allocated to different OFDM subcarriers, added by a cyclic prefix (CP) to mitigate inter-symbol interference (ISI) caused by multi-path fading. The data transmitted to the $k$-th user of the $m$-th subcarrier, $m=1,2,\cdots,M$, is referred to as ${x}_{m,k}$.
Due to the sparse scattering environment of the mmWave frequency [18], we adopt the Saleh-Valenzuela channel Model for the multi-path channel $\mathbf{h}_{m,k}\in\mathbb{C}^{N_{t}\times 1}$ between the BS and the user $k$ at the $m$-th subcarrier as
$\mathbf{h}_{m,k}=\sum_{l=1}^{L_{k}}\beta_{m,k,l}e^{-j2\pi f_{m}\tau_{k,l}}\mathbf{a}_{m}(\theta_{k,l},\phi_{k,l}),\vskip-2.84526pt$ (1)
@@
$\vskip-2.84526pt\mathbf{a}_{m}(\theta,\phi)$ $=\frac{1}{\sqrt{N_{t}}}[1,\cdots,e^{jk_{m}d(n_{v}\sin\phi\sin\theta+n_{h}\cos\theta)},\cdots,$
$e^{jk_{m}d((N_{v}-1)\sin\phi\sin\theta+(N_{h}-1)\cos\theta)}]^{T},\vskip-2.84526pt$ (2)
@@
Building upon the aforementioned description of the MIMO-OFDM transceiver, we select critical yet challenging tasks including channel estimation, MIMO detection, channel decoding, and multi-user precoding, while ingoring the relatively trivial components (e.g., channel encoding, which requires only a matrix multiplication operation). The task selection aims to establish a unified FM capable of addressing the most demanding aspects of a transceiver. Furthermore, we incorporate user localization as an essential task, given its fundamental role in 6G ISAC systems, a pivotal direction for next-generation wireless networks. The designed multi-task FM aims to handle tasks across various physical layers as comprehensively as possible. In the next subsection, we will formulate the selected tasks individually.
@@ III Multi-task Environment-aware Foundation Model (MUSE-FM)
Figure 3: Vision-transformer based scene encoder.
@@ III-A Overall Framework
The framework of the proposed model is presented in Fig. 2, which includes the scene encoder, prompted-guided unified encoder-decoder pair, and FM backbone. To address scenario limitation, leveraging the strong correlation between physical environments and channel characteristics [20], we attempt to incorporate environmental information to allow cross-scenario generalization in multi-task FM. Specifically, the spatial structure of the physical environment, including the area building distribution as well as the BS position, determines the propagation path information of the wireless channel, and thus directly impacts channel characteristics. Therefore, involving the environmental knowledge as a multi-modal input provides a promising way to inform FMs of the channel distribution and improves the generalization ability in different scenarios. We implement it through a Vision Transformer-based encoder that extracts environmental and channel prior knowledge from scene graphs. The task-agnostic prior knowledge is subsequently fused with wireless data to enhance the model’s environmental adaptability. The specific design of the scene encoder is described in Section III-B.
@@
To handle data with heterogeneous formats and distributions across multiple tasks, we propose a prompt-guided unified data encoder whose parameters are dynamically generated by task-specific instructions through hypernetworks. In particular, the task instruction is fed to a hypernetwork to dynamically generate parameters, which is utilized in the data encoder. This architecture allows the instruction to “guide” the encoder to adaptively extract task-specific features according to each task’s objectives and data characteristics. To achieve this, two key components are included: a parameter generator, i.e., which is used to generate task-specific parameters from task instruction via a text encoder; a prompted-guided unified encoder-decoder pair, to utilize the generated parameters to project the data of multiple tasks into the shared FM space (or generate the final outputs in multiple tasks). The two components are elaborated in Section III-C and III-D.
@@
The token embeddings from scene graphs and wireless data (e.g., CSI and received symbols) are concatenated to form the input embeddings for the FM backbone. The FM backbone consists of position embedding and stacked classical transformer blocks, which is illustrated in Section III-E. Equipped with large-scale parameters and a layered structure, the FM backbone exhibits expressive learning capacity for multiple tasks. Besides, atthention mechanism enables effective fusion of multi-modal inputs. It assimilates related channel features from task-agnostic prior knowledge for different tasks and integrates it with wireless data, ensuring cross-scenario feature extraction. Finally, similar to the data encoder, equipped with task-aware parameters, the prompt-guided unified decoder generates the output for different tasks.
MUSE-FM is designed to handle a wide range of physical-layer tasks. When encountering novel tasks, its uniform architecture allows for extension through minimal task-specific adaptation by designing a dedicated task instruction and fine-tuning the model with limited data samples.
@@ III-B Scene Encoder
The scene encoder aims to extract task-agnostic priors of channel and environment from scene graphs. The structure of the scene encoder to effectively exploit environmental information is illustrated in Fig. 3. Direct utilization of raw 3D environmental data is challenging due to its high dimensionality. We therefore implement preprocessing, transforming environmental information into compact representations suitable for encoder input. Given that top-down views holistically encapsulate the spatial structure of the environment, that is, scene layouts and obstacle distributions, we extract it to construct the scene graph [21]. This scene graph is then discretized into a matrix form, where the value 1 represents the obstacles, and the value 0 represents free space.
@@ III-D1 Pre-processor and unified data encoder
In Fig. 4, we take channel estimation as an example to illustrate how the original data input $\mathrm{X}_{n}^{\rm pre}$ is preprocessed and handled by the unified data encoder, obtaining $\mathrm{X}_{n}^{\rm emb}$.
Finnaly, the data embeddings $\mathrm{X}_{n}^{\rm emb}$ are concatenated with the scene embeddings $\mathrm{X}_{s}^{\rm emb}$. Besides, we use the standard approach of adding an extra learnable “classification token” (cls token) to the sequence, to form the final input of the FM backbone. We refer to the input as $\mathrm{X}_{\rm FM}^{\rm emb}$.